### v1

# Extraction of ground-state nuclear deformations from ultra-relativistic heavy-ion collisions: Nuclear structure physics context

Abstract
The collective-flow-assisted nuclear shape-imaging method in ultra-relativistic heavy-ion collisions has recently been used to characterize nuclear collective states. In this paper, we assess the foundations of the shape-imaging technique employed in these studies. We conclude that, on the whole, the discussion regarding low-energy nuclear physics is confusing and the suggested impact on nuclear structure research is overstated. Conversely, efforts to incorporate existing knowledge on nuclear shapes into analysis pipelines can be beneficial for benchmarking tools and calibrating models used to extract information from ultra-relativistic heavy ion experiments.
I Introduction
The shapes of nuclei colliding in ultra-relativistic heavy-ion collisions (UHIC) influence the geometry of the quark-gluon plasma (QGP) created which, in turn, affects the momentum distribution of the particles produced [1, 2]. The premise of the UHIC studies is rooted in the fact that the resulting anisotropic (hydrodynamic) expansion of the QGP [3] converts the initial spatial asymmetries into momentum anisotropies of the measured particles. A typical method of analyzing this flow consists of characterizing the azimuthal anisotropy by the Fourier coefficients of the flow; the elliptic flow, , is sensitive to the ellipticity of the QGP, while the triangular one, , is sensitive to the triangularity, and so on.
By relating the flow anisotropies to the geometry of the QGP initial state, a shape-imaging method based on UHIC experiments at the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC) has been suggested as a tool to study the geometrical shapes of colliding nuclei [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. In the present study, we look deeper into the assumptions behind the proposed shape-imaging technique. In particular, we critically assess the claim in Ref. [6] that this approach “not only refines our understanding of the initial stages of the collisions but also turns high-energy nuclear experiments into a precision tool for imaging nuclear structures.”
The paper is organized as follows. Section II discusses the concept of snapshots behind the shape-imaging method, and Sec. III covers the notion of “shape fluctuations.” In Sec. IV we recall the vast body of available knowledge on nuclear shapes from low-energy measurements. The questions related to the uncertainty quantification (UQ) of the shape-imaging method are raised in Sec. V. The conclusions of the paper are given in Sec. VI.
II Symmetries, time scales, and snapshots
Quantum mechanics frequently defies simple intuition. At the same time, simple, intuitive models based on experience often help us to understand the world. These two points should not be confused. Recently, several reports claiming a rigorous determination of nuclear shapes from UHIC experiments have done so to a greater or lesser extent [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
A fundamental theorem of quantum mechanics states that an eigenstate of a symmetry-conserving Hamiltonian must belong to a representation of the symmetry group. Therefore, the ground states of 238U and 20Ne, which a rotationally-invariant nuclear Hamiltonian governs, belong to the representation, i.e., they are invariant under spatial orientation and reflections. In other words, they are perfectly spherical states in the beam’s reference frame (or perfectly axial in the laboratory frame due to the huge relativistic contraction). Similarly, any ground or excited state with [20, 21] for which the laboratory quadrupole moment is non-zero is covariant under spatial orientation (its rotation mixes the magnetic substates in a specific way). We focus on the case for the following discussion, but note that all arguments apply to any of the others.
What is the shape, then, of the 238U nucleus in its 0+ ground state and how can it be measured? The answer to these questions is irrespective of whether we use UHIC, electron scattering, or Coulomb excitation. A concept that addresses this properly is spontaneous symmetry breaking – an essential notion ubiquitous across all the various domains of physics. As exposed by Anderson in his seminal article [22], it encompasses finite systems (such as molecules or nuclei), very large systems (like crystals), and infinite systems (such as quantum fields). It does not contradict the aforementioned fundamental theorem of quantum mechanics; on the contrary, it allows us to understand what it means to measure the shape of a quantum object.
A false friend in this understanding is an intuition by which one may picture the deformed 238U nucleus as a rotating and vibrating classical object with corresponding time scales inferred from the energies of the quantum rotational and vibrational states [4]. This is a misleading picture – the ground state of 238U studied in the experiments is a stationary eigenstate, it does not rotate or vibrate, it does not change with time, and, thus, no instantaneous “snapshot” [11] of the shape can be taken, no matter the time scale of the experiment performed.
A proper quantum-mechanical intuition that helps us to understand the shape of the 238U 0+ ground state is a model of a spontaneously symmetry-broken deformed wave function. This is not the wave function that reaches the detector in the UHIC experiment. However, this deformed wave function can be used to build an intuitive model of the 238U 0+ ground state by constructing a linear combination of wave functions differently oriented in space and having different shapes. This model of the 0+ ground state is still time-independent; it does not rotate or vibrate. Most importantly, one must remember that in this model state the wave functions of different orientations and shapes are mixed, not their squares or density distributions, as pictured, e.g., in the tip-tip or body-body configuration schematics of Ref. [5].
As is the case for any other quantum-mechanical wave function, a single experiment, when properly designed, may project the wave function on a given value of its “position” (its orientation and shape). However, a series of several identical experiments will give different values of positions. In particular, for the 0+ 238U ground state, all the different possible orientations will come out with equal probabilities, and the information on the shape will be inaccessible. Indeed, it must be so, as the shape of the 0+ state is spherical.
A similar discussion can be made about the “bowling pin” schematics of Refs. [16, 17] symbolizing a pear-shaped nucleus, 20Ne. Indeed, in the laboratory system, the ground-state wave function of 20Ne is perfectly spherical and reflection-symmetric, i.e., it does not point in a specific direction.
So, how then can we measure the shape of a spontaneously symmetry-broken state of 238U or 20Ne? We cannot. The reason is simple: such a stationary state does not exist in nature and, therefore, is unavailable for experimentation. In quantum mechanical modeling, the deformed symmetry-broken state is a wave packet, a coherent superposition of states with different angular momenta. A possible experimental situation where such a state might briefly appear is in a heavy-ion fusion-evaporation reaction, where two ions fuse and rotate together at very high angular momentum. However, immediately after the first de-excitation photon is detected in a -ray detector, the wave packet collapses into a high-angular-momentum eigenstate, which then emits a series of photons and rapidly relaxes to the ground state. Such states are not the ones probed in the UHIC experiments, however.
At this point, we are ready to discuss what is measured in the UHIC experiments. The key point is that one does not measure nuclear moments in these events. Instead, through event reconstruction, the positions of individual nucleons are identified based on measurements of outgoing particles. Multiple identical collisions then provide access to different quantum probability distributions. In short, every UHIC takes a snapshot of the spatial positions of the nucleons, not a snapshot of nuclear rotations or vibrations [11]. In analogy, this can be related to Coulomb explosion experiments, which yield geometrical images of individual molecules [23, 24], see also Ref. [4].
Any measurement of the probability distribution to find one nucleon at any given position in space maps the (exact) one-body density of the nucleus. It is equal to the modulus squared of the (exact) many-body wave function, which depends on the coordinates of particles, integrated over coordinates. The one-body density of any 0+ state is spherical. To determine the shape, one must use the concept of conditional probability, that is, the probability of finding one nucleon at a specific position in space under the condition that another nucleon occupies another position in space. Such a conditional probability is related to the (exact) two-body density (the integral of the exact wave function’s modulus squared over coordinates). It has an axial shape because it must be symmetric with respect to the line connecting the two spatial points in question. In this way, the exact axial shape of a 0+ state can be inferred. Therefore, if the UHIC measurements can extract such conditional probabilities, the axial deformation can be assessed, also in a 0+ state.
Examples of such theoretical analyses, involving two-particle correlations, can be found in, e.g., Refs. [25, 26], where the mean-field, symmetry-breaking, and symmetry-restored solutions are compared with the exact solutions of the problem, with many aspects of the arguments mentioned above clearly and explicitly presented. See also examples of calculations in Refs. [27, 16] of correlated samplings of nucleon positions. In the context of the present discussion, it is useful to note that a local approximation to the two-particle correlation function, based on the density-matrix expansion of the two-body density, is the nucleonic localization function [28], which was originally introduced to visualize bond structures in molecules. This nucleonic localization function was employed in mean-field calculations to study cluster configurations in nuclei [29, 30], cluster structures in fission [31], formation of clusters in both heavy-ion reactions [32] and high-spin states [33].
In the same way, to calculate or measure triaxial deformations, one must consider the (exact) three-body density or measure correlations between positions of three nucleons. This is not what was done in Ref. [12], in which the triaxiality of 129 Xe was inferred based on LHC data.
In summary, ultra-relativistic collisions take place in the laboratory frame. Since 238U is spherical in its ground state and angular momentum is conserved, a deformed collision image is impossible, based on physics principles, regardless of the collision mechanism. (Likewise, a direct laboratory-system measurement of the electric dipole moment of the ammonia molecule in its ground state is impossible as the molecule has good parity [22].)
In other words, in the beam’s coordinate system, the heavy ion impinging on target is described by a rotationally-invariant many-body wave function that depends on the coordinates of the nucleons, not just on the nucleon densities oriented in space. It is following the collision event that the wave function collapses into one symmetry-breaking component, characterized by the nucleons’ positions (planar image of the wave function in the coordinate representation). Hence, to extract the initial QGP configuration, a simple one-body picture of the deformed nuclear density is insufficient.
III Static shapes and dynamic shape fluctuations
Publications by the STAR collaboration [4, 5] introduce the notion of “instantaneous shapes” and “long-timescale quantum fluctuations, making direct observation challenging.” It is claimed that “Nuclear shapes, even in ground states, are not fixed. They exhibit zero-point quantum fluctuations involving various collective and nucleonic degrees of freedom at different timescales. These fluctuations superimpose on each other in the laboratory frame” [4].
We wish to point out that there are no quantum fluctuations in any eigenstate of any quantum system. The shape of a projectile/target nucleus entering the collision does not fluctuate; there is no time-dependence involved, and, thus, there is no timescale. One should not confuse the model of a collective rotational or vibrational state, described as a time-dependent Slater determinant, with the reality of a quantum state that is accessible in an experiment. Similarly, one should not confuse dynamical fluctuations with the quantum uncertainty of an observable. An observable, the quantum quadrupole moment operator, defines the nuclear shape. The nuclear ground state is not an eigenstate of the quadrupole operator; therefore, any measurement of the quadrupole operator in the nuclear ground state must lead to the standard quantum dispersion of the results. Such dispersion is not related to any collective motion.
While a theoretical jargon exists pertaining to static deformations (related to symmetry-breaking deformed mean-field solutions) and dynamical zero-point effects, or fluctuations (representing the dynamical beyond-mean-field corrections), the corresponding many-body wave functions are always stationary. Likewise, the jargon referring to the “intrinsic” reference frame is unhelpful in this context; it is better to use the proper quantum-mechanical notion of a symmetry-broken state.
IV Nuclear deformations: a concept not entirely unknown
Nuclear deformations are parameters characterizing the anisotropy of the nuclear shape. These characteristics are not fundamental nuclear properties, as they are deduced from various measured observables in a model-dependent way [34]. In low-energy nuclear physics, nuclear shapes have been characterized for decades using observables, such as quadrupole moments, obtained from a wide variety of experimental techniques. Since the seminal works by Schüler and Schmidt [35] and Casimir [36, 37], nuclear electric quadrupole moments have been extracted, evaluated, and tabulated for nuclear ground states or excited states with total angular momenta of , from a number of high-precision measurements [38, 39], often quoted with uncertainties of 1% or less for stable nuclei. The nuclear quadrupole moment measures the deviation of the nuclear charge distribution from sphericity in a given nuclear state. In fact, Ref. [38] lists no less than 50 experimental techniques that form the basis for the moments compiled. These range from the hyperfine splitting in mesic atoms [40, 41, 42], to a variety of laser spectroscopic techniques that are even applicable to short-lived rare isotopes [43, 44], to scattering approaches such as electron scattering [45, 46, 47] or heavy-ion-induced Coulomb excitation [48, 49]. For a recent review on the history of nuclear shapes and experimental methods, the reader is referred to the recent review [50].
This body of precision data from low-energy nuclear physics experiments seems largely unknown to the UHIC community and its existence is in direct contradiction with statements such as found in Ref. [12]: “Additionally, for odd-mass and odd-odd nuclei, quantum mechanics entangles the contributions from collective intrinsic deformation and individual single-particle states to the quadrupole moment, such that the former cannot be uniquely determined from a spectroscopic experiment.” One example from Ref. [38] is the moment, b, quoted for the ground state of the odd-mass 197Au nucleus from muonic x-ray spectroscopy, a nucleus that was much discussed by the UHIC community, e.g., in Refs. [13, 4]. Additionally, taking any odd-odd nucleus, here 28Na, from Ref. [38] reveals that moments can be uniquely determined from spectroscopic experiments, e.g., for 28Na from -NMR spectroscopy where a high degree of nuclear polarization was achieved via optical pumping with a collinear laser beam [51]. We note that this was accomplished for a short-lived rare isotope which cannot be accessed by UHIC at RHIC or the LHC.
For most stable and many short-lived nuclei, including for odd-mass and even-even ones, which have a zero ground-state total angular momentum, a vast body of low-energy data exists from Coulomb excitation [48, 49] as well as from electron scattering [52, 53], sometimes even from Mössbauer spectroscopy [54], on the deformation of excited states, including for quadrupole [55, 56] and higher-order deformations such as the octupole (see Refs. [57, 58]) and hexadecapole ones (e.g., [59, 60, 61, 62]). When appropriate in the scientific context, the results from such low-energy measurements use the , , and [63, 48] parameterizations of the deformation as adopted by the UHIC community. Often, the direct use of reduced electromagnetic transitions strengths, , , and proved to be more insightful and scientifically appropriate in the context of the field. The case has yet to be made for a unique nuclear structure insight from UHIC on 238U [4, 14, 64], 20Ne [16, 17], 96Zr and 96Ru [15], 129Xe [12], 197Au [13], and 150Nd [18], given the substantial and quantum-mechanically correctly interpreted body of low-energy data on nuclear deformation [55, 38, 56].
Theoretically, nuclear deformations can be obtained from calculated one-body nuclear densities of symmetry-broken states, including charge density, proton density, or neutron density. A substantial body of literature exists on local and global predictions of nuclear shape deformations, primarily obtained with various flavors of mean-field theory and its extensions. Global surveys can be found in, e.g., Refs. [65, 66, 67, 68, 69, 70, 71, 72, 73] and the theoretical databases MassExplorer [74] and BrusLib [75]. As discussed in Ref. [34], there is a weak model dependence of nuclear deformations of well-deformed nuclei as they reflect the geometries of the valence Hartree-Fock (Kohn-Sham, canonical, Nilsson) orbits. While ab-initio computations have been advocated in several UHIC imaging papers [6, 16], the current state of the ab initio description of deformations is either not satisfactory [76] or in its infancy [77], and more work is needed to make it globally competitive with well-calibrated mean-field-based methods.
V Uncertainty quantification
Uncertainty quantification is at the heart of any pipeline connecting models that span disparate scales. By quantifying the models’ discrepancies and deficiencies, one can construct a rigorous statistical framework that provides a means to transfer information both forwards and backwards within this multiscale (or multifidelity) modeling framework. In the context of the current work, the forward direction involves propagating uncertainties from quantified nuclear models through to final-state observables in UHIC. The natural next step is then to have information flowing in the other direction in the pipeline. This provides an opportunity to solve the inverse problem and perform statistical inference on shapes and other correlated observables accessible within the chosen nuclear structure model. Both steps represent exciting developments and are excellent examples of the confluence of advances in theoretical modeling, experimental analyses, and computational statistics and data science in modern nuclear physics.
A caveat emerges, however, when one attempts to enact this flow of information for complex multiscale models. There has been considerable progress in quantifying the uncertainty in the context of UHIC in recent years: see Refs. [83, 84, 5] for a non-exhaustive selection of works with a detailed discussion of UQ practices. However, without careful consideration of the full uncertainties for each component of the entire multistage process, one runs the risk of misspecified uncertainties polluting parts of the pipeline. Fits and calibrations, in general, will lead to distributions of tunable parameters that allow one, within a given credible interval, to explain an equivalent subset of available data. If those tuneable parameters control a model with an intrinsic model deficiency or inability to describe different classes of observables simultaneously, the interpretation of those model outputs and parameters must be viewed with skepticism.
This fact is particularly concerning if one wishes to use information from UHIC alone to constrain nuclear structure models or to provide information on nuclear shapes. In the simplest applications of this framework, the nuclear structure model has been chosen as a deformed Fermi form factor with deformation parameters constraining the bulk shape of the nucleus. In this example, the calibration of the model involves exploring the various parameters that define the shape and determining a statistical ensemble of shapes that best agree with the downstream UHIC data. Without a simultaneous constraint or consideration of binding energies, charge radii, or deformations from other observables, it is difficult to know the full extent of the discovery potential of such analyses.
It is also essential to acknowledge the fact that any conclusion drawn in such an analysis is inherently model-dependent. The experimental data, on an event-by-event basis, undoubtedly encode intricate details about the correlated nuclear system. Decoding this information, however, necessitates a sophisticated theoretical modeling framework that translates UHIC observations into quantitative measures of nuclear deformation. This framework typically involves several interconnected stages, each with its own set of assumptions and uncertainties.
Centrality, for instance, in heavy-ion collisions is typically determined by measuring the number of charged particles produced in the collision and then comparing this number to a Monte Carlo Glauber calculation or another similar framework [85, 86, 87, 88]. This approach, while widely used, has certain limitations. The accuracy of centrality determination relies on the validity of the chosen model, which has inherent assumptions about the nucleon-nucleon interactions and the density distribution of nucleons within the nucleus. Additionally, the efficiency of detecting charged particles and the methods used to distinguish between particles originating from the collision and those from secondary decays can introduce uncertainties. The subsequent hydrodynamic evolution of the QGP is another example as it relies on equations of state and transport coefficients that are not fully constrained by experimental data, and even then, there are multiple competing models to propagate the plasma after formation [89]. Finally, the process of hadronization, whereby the QGP converts back into hadrons, introduces further uncertainties related to the fragmentation, clusterization, and subsequent decay of these particles [90, 91].
The complexity and non-uniqueness of this modeling framework raise concerns regarding the robustness and reproducibility of results drawn from the method. Different research groups, employing subtly different implementations of the same underlying stages, can arrive at different conclusions regarding the extracted deformations of the colliding system. These discrepancies can arise from variations in the specific data sets considered in the fit, the choice of initial conditions, the details of the hydrodynamic evolution, the treatment of hadronization, and other factors.
Finally, it is worth discussing the required properties and fidelity of the nuclear structure models to be useful in this pipeline. At a minimum, one should consider ‘reasonable’ models of nuclei. A convenient definition in this case would be any model that allows one to simultaneously describe (and predict) observables in low-energy nuclear structure to a reasonable accuracy. For the forward direction in the pipeline, this would also ideally involve quantified models that have been calibrated to a wide class of nuclear observables and have been validated on the class of observables that are of most interest in any particular study. The uncertainties of the initial state should then be propagated through the pipeline, from the hydro phase of the QGP to the hadronization phase. This already provides an excellent anchor point for UHIC studies to ensure that the nuclear structure input is well validated, providing further confidence in any UQ efforts for the UHIC simulation pipeline itself.
For the reverse direction, the core question is “are the constraints and resulting uncertainties sufficient to distinguish between models used to predict the initial state?”. If the discussed pipeline is rather coarsely constraining, then one should only expect differences to arise from large-scale, bulk changes in the nucleus. For models meeting the reasonableness criteria above, these bulk properties are likely to be very similar. In particular, employing ab-initio nuclear structure models for inference on nuclear properties of any sort requires quantified Hamiltonians and detailed model discrepancy studies to validate those models on nuclear deformations and to render any conclusions statistically significant. While this discussion has focused on nuclear shapes, this is a core point for the extraction of any properties of nuclei. For example, the investigation in Ref. [18] suggests UHIC experiments as a novel probe of matrix elements central to the search for neutrinoless double-beta decay, although without considering the constraining potential of those UHIC observables in the context of other available nuclear structure observables (see, e.g., Ref . [92] for a multitude of factors impacting the matrix element for neutrinoless double-beta decay). More fundamentally, one needs to demonstrate the suitability of the structure model itself for such studies by showing that, when considering data from other sources, the results are not in conflict with well-known nuclear properties.
VI Conclusions
“Extraordinary claims require extraordinary evidence” [93]. The purpose of this paper is to shed light on the collective-flow-assisted nuclear shape imaging employed in the analysis of data from UHIC, and to assess the claims of high relevance of these experiments to nuclear structure research. While there is little doubt that the proper treatment of the geometries of colliding heavy ions is essential for the characterization of the initial QGP configuration and, hence, the extraction of fundamental properties of QGP from UHIC, we believe that the proposed nuclear shape-imaging method is based on several flawed assumptions and, hence, its usefulness as a “discovery tool for exploring nuclear structure” [4] has been overstated.
First, the modeling of the initial QGP state formed following the heavy-ion collision involves the description of the relative position identification of the individual nucleons. To this end, the information contained in the one-body nucleonic density, as presented in the majority of papers dealing with nuclear imaging from UHIC, is insufficient. The proper tool for this imaging of the nucleon’s position distributions is the two-body density (to map axial distributions) or the three-body one (to map triaxial shapes). The schematics depicting the deformed orientations of nuclei are misleading, as even-even nuclei, such as 238U or 20Ne, have isotropic ground-state wave functions in the laboratory system.
Second, the notions of “instantaneous shapes,” “long-timescale quantum fluctuations,” and “zero-point quantum fluctuations involving various collective and nucleonic degrees of freedom at different timescales” [4] make little physical sense as the many-body wave function of the colliding ion is stationary.
Third, rich databases exist of shape deformations measured in low-energy experiments, as well as of theoretical deformations computed within nuclear mean-field theory. The data on stable nuclei are particularly rich and seem more than sufficient to inform models of the initial QGP state.
Fourth, the existing precise information on nuclear shapes should be utilized by nuclear UHIC imaging practitioners to validate analysis techniques and control model uncertainties and discrepancies. Currently, a comprehensive and combined uncertainty quantification analysis has not been carried out.
In summary, we find that the impact of nuclear UHIC imaging on low-energy nuclear structure is still limited. Nuclear shape parameters for stable or long-lived nuclei are well understood, having been studied extensively over many years using a wide range of techniques, both direct and indirect. Given the significant uncertainties associated with the UHIC imaging process, which relies on phenomenological models, the accuracy of current systematic nuclear structure measurements and calculations appears sufficient to inform the modeling of the QGP initial state.
Acknowledgements.
This material is based upon work supported by the U.S. Department of Energy under Award Numbers DE-SC0013365 and DE-SC0023633 (Office of Science, Office of Nuclear Physics), DE-SC0023175 (Office of Science, NUCLEI SciDAC-5 collaboration), DE-FG02-97ER41041, DE-FG02-97ER41033, and the National Science Foundation under award number 2004601 (CSSI program, BAND collaboration). This work was also partially supported by the STFC Grant Nos. ST/P003885/1 and ST/V001035/1.References
- Shuryak [2017] E. Shuryak, Strongly coupled quark-gluon plasma in heavy ion collisions, Rev. Mod. Phys. 89, 035001 (2017).
- Busza et al. [2018] W. Busza, K. Rajagopal, and W. van der Schee, Heavy ion collisions: The big picture and the big questions, Ann. Rev.Nucl. Part. Sci. 68, 339–376 (2018).
- Heinz and Schenke [2024] U. Heinz and B. Schenke, Hydrodynamic description of the quark-gluon plasma (2024), arXiv:2412.19393 [nucl-th] .
- Abdulhamid et al. [2024] M. I. Abdulhamid et al., Imaging shapes of atomic nuclei in high-energy nuclear collisions, Nature 635, 67–72 (2024).
- The STAR Collaboration [2025] The STAR Collaboration, Imaging nuclear shape through anisotropic and radial flow in high-energy heavy-ion collisions (2025), arXiv:2506.17785 [nucl-ex] .
- Jia et al. [2024] J. Jia et al., Imaging the initial condition of heavy-ion collisions and nuclear structure across the nuclide chart, Nucl. Sci. Tech. 35, 220 (2024).
- Giacalone [2020a] G. Giacalone, Constraining the quadrupole deformation of atomic nuclei with relativistic nuclear collisions, Phys. Rev. C 102, 024901 (2020a).
- Giacalone [2020b] G. Giacalone, Observing the deformation of nuclei with relativistic nuclear collisions, Phys. Rev. Lett. 124, 202301 (2020b).
- Giacalone et al. [2021a] G. Giacalone, J. Jia, and V. Somà, Accessing the shape of atomic nuclei with relativistic collisions of isobars, Phys. Rev. C 104, L041903 (2021a).
- Giacalone et al. [2021b] G. Giacalone, J. Jia, and C. Zhang, Impact of nuclear deformation on relativistic heavy-ion collisions: Assessing consistency in nuclear physics across energy scales, Phys. Rev. Lett. 127, 242301 (2021b).
- Jia [2022] J. Jia, Shape of atomic nuclei in heavy ion collisions, Phys. Rev. C 105, 014905 (2022).
- Bally et al. [2022] B. Bally, M. Bender, G. Giacalone, and V. Somà, Evidence of the triaxial structure of at the large hadron collider, Phys. Rev. Lett. 128, 082301 (2022).
- Bally et al. [2023] B. Bally, G. Giacalone, and M. Bender, The shape of gold, Eur. Phys. J. A 59, 58 (2023).
- Ryssens et al. [2023] W. Ryssens, G. Giacalone, B. Schenke, and C. Shen, Evidence of hexadecapole deformation in uranium-238 at the relativistic heavy ion collider, Phys. Rev. Lett. 130, 212302 (2023).
- Zhang and Jia [2022] C. Zhang and J. Jia, Evidence of quadrupole and octupole deformations in and collisions at ultrarelativistic energies, Phys. Rev. Lett. 128, 022301 (2022).
- Giacalone et al. [2024] G. Giacalone, B. Bally, G. Nijs, S. Shen, T. Duguet, J.-P. Ebran, S. Elhatisari, M. Frosini, T. A. Lähde, D. Lee, B.-N. Lu, Y.-Z. Ma, U.-G. Meißner, J. Noronha-Hostler, C. Plumberg, T. R. Rodríguez, R. Roth, W. van der Schee, and V. Somà, The unexpected uses of a bowling pin: exploiting 20Ne isotopes for precision characterizations of collectivity in small systems (2024), arXiv:2402.05995 [nucl-th] .
- Giacalone et al. [2025] G. Giacalone, W. Zhao, B. Bally, S. Shen, T. Duguet, J.-P. Ebran, S. Elhatisari, M. Frosini, T. A. Lähde, D. Lee, B.-N. Lu, Y.-Z. Ma, U.-G. Meißner, G. Nijs, J. Noronha-Hostler, C. Plumberg, T. R. Rodríguez, R. Roth, W. van der Schee, B. Schenke, C. Shen, and V. Somà, Anisotropic flow in fixed-target Ne collisions as a probe of quark-gluon plasma, Phys. Rev. Lett. 134, 082301 (2025).
- Li et al. [2025] Y. Li, X. Zhang, G. Giacalone, and J. Yao, Benchmarking nuclear matrix elements of decay with high-energy nuclear collisions (2025), arXiv:2502.08027 [nucl-th] .
- Fortier et al. [2025] N. M. Fortier, S. Jeon, and C. Gale, Heavy-ion collisions as probes of nuclear structure, Phys. Rev. C 111, L011901 (2025).
- Bragg [1948] J. K. Bragg, The interaction of nuclear electric quadrupole moments with molecular rotation in asymmetric-top molecules. I, Phys. Rev. 74, 533–538 (1948).
- Bohr [1951] A. Bohr, On the quantization of angular momenta in heavy nuclei, Phys. Rev. 81, 134–138 (1951).
- Anderson [1972] P. W. Anderson, More is different. Broken symmetry and the nature of the hierarchical structure of science, Science 177, 393 (1972).
- Vager et al. [1989] Z. Vager, R. Naaman, and E. Kanter, Coulomb explosion imaging of small molecules, Science 244, 426 (1989).
- Miller [2022] J. L. Miller, Coulomb-explosion imaging tackles an 11-atom molecule, Phys. Today 75, 12–14 (2022).
- Romanovsky et al. [2006] I. Romanovsky, C. Yannouleas, L. O. Baksmaty, and U. Landman, Bosonic molecules in rotating traps, Phys. Rev. Lett. 97, 090401 (2006).
- Sheikh et al. [2021] J. A. Sheikh, J. Dobaczewski, P. Ring, L. M. Robledo, and C. Yannouleas, Symmetry restoration in mean-field approaches, J. Phys. G 48, 123001 (2021).
- Shen et al. [2023] S. Shen, S. Elhatisari, T. A. Lähde, D. Lee, B.-N. Lu, and U.-G. Meißner, Emergent geometry and duality in the carbon nucleus, Nat. Commun. 14, 2777 (2023).
- Becke and Edgecombe [1990] A. D. Becke and K. E. Edgecombe, A simple measure of electron localization in atomic and molecular systems, J. Chem. Phys. 92, 5397–5403 (1990).
- Reinhard et al. [2011] P.-G. Reinhard, J. A. Maruhn, A. S. Umar, and V. E. Oberacker, Localization in light nuclei, Phys. Rev. C 83, 034312 (2011).
- Nakatsukasa and Hinohara [2023] T. Nakatsukasa and N. Hinohara, Local -removal strength in the mean-field approximation, Phys. Rev. C 108, 014318 (2023).
- Zhang et al. [2016] C. L. Zhang, B. Schuetrumpf, and W. Nazarewicz, Nucleon localization and fragment formation in nuclear fission, Phys. Rev. C 94, 064323 (2016).
- Schuetrumpf and Nazarewicz [2017] B. Schuetrumpf and W. Nazarewicz, Cluster formation in precompound nuclei in the time-dependent framework, Phys. Rev. C 96, 064608 (2017).
- Li et al. [2020] T. Li, M. Z. Chen, C. L. Zhang, W. Nazarewicz, and M. Kortelainen, Nucleon localization function in rotating nuclei, Phys. Rev. C 102, 044305 (2020).
- Nazarewicz and Ragnarsson [1996] W. Nazarewicz and I. Ragnarsson, Nuclear Deformations, in Handbook of Nuclear Properties, edited by D. Poenaru and W. Greiner (Clarendon Press, 1996) p. 80.
- Schüler and Schmidt [1935] H. Schüler and T. Schmidt, Über Abweichungen des Atomkerns von der Kugelsymmetrie, Z. Phys. 94, 457–468 (1935).
- Casimir [1935] H. Casimir, Über die Hyperfeinstruktur des Europiums, Physica 2, 719–723 (1935).
- Casimir [1936] G. B. H. Casimir, On the Interaction between Atomic Nuclei and Electrons (Teylers Tweede Genootschap, Haarlem, 1936).
- N. J. Stone [2021] N. J. Stone, INDC International Nuclear Data Committee - Table of nuclear electric quadrupole moments, https://www-nds.iaea.org/publications/indc/indc-nds-0833.pdf (2021), accessed: 29-06-2025.
- Pyykkö [2018] P. Pyykkö, Year-2017 nuclear quadrupole moments, Mol. Phys. 116, 1328–1338 (2018).
- Fitch and Rainwater [1953] V. L. Fitch and J. Rainwater, Studies of x-rays from mu-mesonic atoms, Phys. Rev. 92, 789–800 (1953).
- Cooper and Henley [1953] L. N. Cooper and E. M. Henley, Mu-mesonic atoms and the electromagnetic radius of the nucleus, Phys. Rev. 92, 801–811 (1953).
- Wheeler [1953] J. A. Wheeler, Mu meson as nuclear probe particle, Phys. Rev. 92, 812–816 (1953).
- Neyens [2003] G. Neyens, Nuclear magnetic and quadrupole moments for nuclear structure research on exotic nuclei, Rep. Prog. Phys. 66, 633 (2003).
- Yang et al. [2023] X. Yang, S. Wang, S. Wilkins, and R. G. Ruiz, Laser spectroscopy for the study of exotic nuclei, Prog. Part. Nucl. Phys. 129, 104005 (2023).
- Hofstadter [1957] R. Hofstadter, Nuclear and nucleon scattering of high-energy electrons, Annu. Rev. of Nucl. Part. Sci. 7, 231–316 (1957).
- Donnelly and Walecka [1975] T. W. Donnelly and J. D. Walecka, Electron scattering and nuclear structure, Ann. Rev. Nucl. Part. Sci. 25, 329–405 (1975).
- Lightbody [1972] J. Lightbody, Electron scattering from one- and two-phonon vibrational states, Phys. Lett. B 38, 475–479 (1972).
- Cline [1986] D. Cline, Nuclear Shapes Studied by Coulomb Excitation, Ann. Rev. Nucl. Part. Sci 36, 683–716 (1986).
- Zielińska [2022] M. Zielińska, Low-energy Coulomb excitation and nuclear deformation, in The Euroschool on Exotic Beams, Vol. VI, edited by S. M. Lenzi and D. Cortina-Gil (Springer International Publishing, Cham, 2022) pp. 43–86.
- Verney [2025] D. Verney, History of the concept of nuclear shape, Eur. Phys. J. A 61, 82 (2025).
- Keim et al. [2000] M. Keim, U. Georg, A. Klein, R. Neugart, M. Neuroth, S. Wilbert, P. Lievens, L. Vermeeren, B. A. Brown, and ISOLDE Collaboration, Measurement of the electric quadrupole moments of 26-29Na, Eur. Phys. J. A 8, 31–40 (2000).
- Cooper et al. [1976] T. Cooper, W. Bertozzi, J. Heisemberg, S. Kowalski, W. Turchinetz, C. Williamson, L. Cardman, S. Fivozinsky, J. Lightbody, and S. Penner, Shapes of deformed nuclei as determined by electron scattering: , , , , , and , Phys. Rev. C 13, 1083–1094 (1976).
- Phan et al. [1989] X. H. Phan, H. G. Andresen, L. S. Cardman, J. M. Cavedon, J. C. Clemens, B. Frois, M. Girod, D. Gogny, D. Goutte, B. Grammaticos, R. Hofmann, M. Huet, P. Leconte, S. K. Platchkov, I. Sick, and S. E. Williamson, Erratum: Electron scattering studies of the ground state rotational band of , Phys. Rev. C 39, 1645–1645 (1989).
- Meeker et al. [1974] R. Meeker, G. Kalvius, B. Dunlap, S. Ruby, and D. Cohen, Hyperfine interactions and nuclear moments in even uranium isotopes from Mössbauer spectroscopy, Nucl. Phys. A 224, 429–436 (1974).
- [55] Evaluated and Compiled Nuclear Structure Data — nndc.bnl.gov, https://www.nndc.bnl.gov/ensdf/, accessed 29-06-2025.
- Pritychenko et al. [2016] B. Pritychenko, M. Birch, B. Singh, and M. Horoi, Tables of E2 transition probabilities from the first 2+ states in even–even nuclei, At. Data Nucl. Data Tables 107, 1–139 (2016).
- Butler and Nazarewicz [1996] P. A. Butler and W. Nazarewicz, Intrinsic reflection asymmetry in atomic nuclei, Rev. Mod. Phys. 68, 349–421 (1996).
- Butler [2020] P. A. Butler, Pear-shaped atomic nuclei, Proc. R. Soc. A 476, 20200202 (2020).
- Eichler et al. [1973] E. Eichler, N. R. Johnson, R. O. Sayer, D. C. Hensley, and L. L. Riedinger, Sign of the hexadecapole moments of and nuclei, Phys. Rev. Lett. 30, 568–571 (1973).
- Ronningen et al. [1977a] R. M. Ronningen, J. H. Hamilton, L. Varnell, J. Lange, A. V. Ramayya, G. Garcia-Bermudez, W. Lourens, L. L. Riedinger, F. K. McGowan, P. H. Stelson, R. L. Robinson, and J. L. C. Ford, and reduced matrix elements of and , Phys. Rev. C 16, 2208–2217 (1977a).
- Ronningen et al. [1977b] R. M. Ronningen, R. B. Piercey, J. H. Hamilton, C. F. Maguire, A. V. Ramayya, H. Kawakami, B. van Nooijen, R. S. Grantham, W. K. Dagenhart, and L. L. Riedinger, Coulomb excitation measurements of reduced and transition matrix elements in , , and , Phys. Rev. C 16, 2218–2222 (1977b).
- Wollersheim and W. Elze [1977] H. Wollersheim and T. W. Elze, Electric quadrupole and hexadecapole moments in the transitional nuclei 150Nd, 152Sm, 154Sm, 154Gd and 156Gd, Nucl. Phys. A 278, 87–94 (1977).
- Wu and Cline [1996] C. Y. Wu and D. Cline, Triaxiality in quadrupole deformed nuclei, Phys. Rev. C 54, 2356–2360 (1996).
- Xu et al. [2024] H.-j. Xu, J. Zhao, and F. Wang, Hexadecapole deformation of from relativistic heavy-ion collisions using a nonlinear response coefficient, Phys. Rev. Lett. 132, 262301 (2024).
- Möller et al. [2008] P. Möller, R. Bengtsson, B. Carlsson, P. Olivius, T. Ichikawa, H. Sagawa, and A. Iwamoto, Axial and reflection asymmetry of the nuclear ground state, At. Data Nucl. Data Tables 94, 758–780 (2008).
- Möller et al. [2016] P. Möller, A. Sierk, T. Ichikawa, and H. Sagawa, Nuclear ground-state masses and deformations: FRDM(2012), At. Data Nucl. Data Tables 109–110, 1–204 (2016).
- S. Goriely et al. [2016] S. Goriely, S. Hilaire, M. Girod, and S. Péru, The Gogny-Hartree-Fock-Bogoliubov nuclear-mass model, Eur. Phys. J. A 52, 202 (2016).
- Erler et al. [2012] J. Erler, N. Birge, M. Kortelainen, W. Nazarewicz, E. Olsen, A. M. Perhac, and M. Stoitsov, The limits of the nuclear landscape, Nature 486, 509–512 (2012).
- Agbemava et al. [2014] S. E. Agbemava, A. V. Afanasjev, D. Ray, and P. Ring, Global performance of covariant energy density functionals: Ground state observables of even-even nuclei and the estimate of theoretical uncertainties, Phys. Rev. C 89, 054320 (2014).
- Cao et al. [2020] Y. Cao, S. E. Agbemava, A. V. Afanasjev, W. Nazarewicz, and E. Olsen, Landscape of pear-shaped even-even nuclei, Phys. Rev. C 102, 024311 (2020).
- Zhang et al. [2022] K. Zhang et al., Nuclear mass table in deformed relativistic Hartree–Bogoliubov theory in continuum, I: Even–even nuclei, At. Data Nucl. Data Tables 144, 101488 (2022).
- Guo et al. [2024] P. Guo et al., Nuclear mass table in deformed relativistic Hartree–Bogoliubov theory in continuum, II: Even-Z nuclei, At. Data Nucl. Data Tables 158, 101661 (2024).
- Zhou and Li [2025] Q. Zhou and Z. Li, Masses and Quadrupole Deformations of Even-Z Nuclei Within a Triaxial Relativistic Hartree–Bogoliubov Model, Particles 8, 10.3390/particles8020057 (2025).
- [74] MassExplorer Database, http://massexplorer.frib.msu.edu/.
- [75] BrusLib Database, http://www.astro.ulb.ac.be/bruslib.
- Vernon et al. [2022] A. R. Vernon, R. F. Garcia Ruiz, T. Miyagi, C. L. Binnersley, J. Billowes, M. L. Bissell, J. Bonnard, T. E. Cocolios, J. Dobaczewski, G. J. Farooq-Smith, K. T. Flanagan, G. Georgiev, W. Gins, R. P. de Groote, R. Heinke, J. D. Holt, J. Hustings, Á. Koszorús, D. Leimbach, K. M. Lynch, G. Neyens, S. R. Stroberg, S. G. Wilkins, X. F. Yang, and D. T. Yordanov, Nuclear moments of indium isotopes reveal abrupt change at magic number 82, Nature 607, 260 (2022).
- Sun et al. [2025] Z. H. Sun, A. Ekström, C. Forssén, G. Hagen, G. R. Jansen, and T. Papenbrock, Multiscale physics of atomic nuclei from first principles, Phys. Rev. X 15, 011028 (2025).
- Faessler [1982] A. Faessler, Description of transitional nuclei, Rep. Prog. Phys. 45, 653 (1982).
- Wood et al. [1992] J. Wood, K. Heyde, W. Nazarewicz, M. Huyse, and P. van Duppen, Coexistence in even-mass nuclei, Phys. Rep. 215, 101–201 (1992).
- Heyde and Wood [2011] K. Heyde and J. L. Wood, Shape coexistence in atomic nuclei, Rev. Mod. Phys. 83, 1467–1521 (2011).
- Reinhard and Otten [1984] P.-G. Reinhard and E. W. Otten, Transition to deformed shapes as a nuclear Jahn-Teller effect, Nucl. Phys. A 420, 173 (1984).
- Nazarewicz [1994] W. Nazarewicz, Microscopic origin of nuclear deformations, Nucl. Phys. A 574, 27 – 49 (1994).
- Nijs et al. [2021] G. Nijs, W. van der Schee, U. Gürsoy, and R. Snellings, Bayesian analysis of heavy ion collisions with the heavy ion computational framework trajectum, Phys. Rev. C 103, 054909 (2021).
- Ehlers et al. [2024] R. Ehlers et al., Measuring jet quenching with a Bayesian inference analysis of hadron and jet data by JETSCAPE, EPJ Web Conf. 296, 15009 (2024).
- ALICE Collaboration [2013] ALICE Collaboration, Centrality determination of pb-pb collisions at tev with ALICE, Phys. Rev. C 88, 044909 (2013).
- Loizides et al. [2015] C. Loizides, J. Nagle, and P. Steinberg, Improved version of the PHOBOS Glauber Monte Carlo, SoftwareX 1-2, 13–18 (2015).
- ALI [2018] Centrality determination in heavy ion collisions, CERN Document Server (2018).
- LHCb collaboration [2022] LHCb collaboration, Centrality determination in heavy-ion collisions with the LHCb detector, J. Instrum 17, P05009 (2022).
- Shen and Yan [2020] C. Shen and L. Yan, Recent development of hydrodynamic modeling in heavy-ion collisions, Nucl. Sci. Tech. 31, 122 (2020).
- Ono [2019] A. Ono, Dynamics of clusters and fragments in heavy-ion collisions, Prog. Part. Nucl. Phys. 105, 139–179 (2019).
- Altmann et al. [2025] J. Altmann, A. Dubla, V. Greco, A. Rossi, and P. Skands, Towards the understanding of heavy quarks hadronization: from leptonic to heavy-ion collisions, Eur. Phys. J. C 85, 1–59 (2025).
- Cirigliano et al. [2022] V. Cirigliano et al., Towards precise and accurate calculations of neutrinoless double-beta decay, J. Phys. G 49, 120502 (2022).
- Sagan [1979] C. Sagan, Broca’s Brain: Reflections on the Romance of Science (Random House, 1979).